OpenStudy (anonymous):

compute the intrgral cos^2(x) * sin2x

OpenStudy (diyadiya):

Is it \[\int\limits \cos^2(x) *\sin2(x)\]

OpenStudy (anonymous):

yes!

OpenStudy (diyadiya):

\[Let \ u=\cos^2(x) \tag1 \]\[du=-2\cos(x)\sin(x) dx\tag2 \] \[\int\limits\limits \cos^2(x)\sin(2x) \ dx= \int\limits\limits \cos^2(x)*2\sin(x) \cos(x) \ dx \]\[Since : \ Sin(2x)=2Sin(x)Cos(x)\] \[\text{Now substituting (1) & (2) }\] \[- \int\limits\limits u \ du\]\[= \frac{-u^2}{2}= \frac{-(\cos^2(x))^2}{2}= \frac{-\cos^4(x)}{2} \]

OpenStudy (anonymous):

Thank you very much!! This answer was extremly helpfull

OpenStudy (diyadiya):

You're Welcome =)